II. Determine the general solution of the given 2nd order linear homogeneous equation. 1. y - 2y + 3y = 0 (ans. y = ci e c
Answer #1

I ol 1. y-24 + 3y =0 Chasa testic eqh 7² - 27 +3=0 2 = +2 + ✓ u-12 2 7= 1 uz i y = C, el cos (2x) tee sin (2) 2 y+y-2y=0D = Izi y = C, Cos 2x & Casinan y (TT) = C. Cos 2 T tc, sinan & since sin2 TI = 0 C = o YLTT) C=0 C Sinza y = y Y(TT) + 2 C